A canteen owner claims that the average meal cost of his usual clients is

Php180. In order to test his own claim, he took a random sample of 30

receipts and computed the mean cost of Php210 with a standard deviation of

Php25. Test the hypothesis at 0.01 level of significance. Show all steps


1
Expert's answer
2022-06-15T17:02:09-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=180"

"H_1:\mu\not=180"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=29" and the critical value for a two-tailed test is "t_c = 2.756386."

The rejection region for this two-tailed test is "R = \{t:|t|> 2.756386\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{210-180}{25/\sqrt{30}}=6.57267"


Since it is observed that "|t|=6.57267> 2.756386=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for two-tailed, "df=29" degrees of freedom, "t=6.57267" is "p=0," and since "p=0<0.01=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is different than 180, at the "\alpha = 0.01" significance level.